x-4√x-5=0
Find x.

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

x-4√x-5=0
---
That is a quadratic whose variable is sqrt(x).
---
Factor::
(sqrt(x)-5)(sqrt(x)+1) = 0
-----
sqrt(x) = 5 or sqrt(x) = -1
-------
x = 25
----
Cheers,
Stan H.
--------------

Answer by Alan3354(69443)

x-4√x-5=0
Find x.
-------------
Sub y for sqrt(x)
--? y^2 - 4y - 5 = 0
(y-5)*(y+1) = 0
y = -1, 5
===============
sqrt(x) = -1
x = 1
----
sqrt(x) = 5
x = 25